OpenStudy (anonymous):

Solve for x: quantity x minus 8 all over 3 equals 12 28 23 20 44

OpenStudy (anonymous):

@bahrom7893 @idvez

OpenStudy (anonymous):

x - 8)/3 = 12 x - 8 = 3 * 12 x - 8 = 36 x = 44